OpenStudy (anonymous):

3a to the power of -3

OpenStudy (anonymous):

@TheOcean

OpenStudy (anonymous):

ok what does a equal to in your question.

OpenStudy (anonymous):

its just a

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

um.

OpenStudy (anonymous):

i have to simplify the equation

OpenStudy (anonymous):

kk. 1 sec

OpenStudy (anonymous):

expression*

jimthompson5910 (jim_thompson5910):

Use the idea that \[\Large x^{-k} = \frac{1}{x^{k}}\]

OpenStudy (anonymous):

umm i dont understand..

jimthompson5910 (jim_thompson5910):

For example \[\Large x^{-2} = \frac{1}{x^2}\] in this case, k = 2

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

So what does \[\Large a^{-3}\] turn into?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

what's the exponent